Complete the following ANOVA summary table using the appropriate formulas.

Source

Sum of Square

df

Mean Square

Fobt

Between groups

1,059

18

Within groups

3,702

167

N/A

Total

4,761

185

N/A

N/A

Calculate the Mean Squared Between (bn)?

Calculate the Mean Squared Within (wn)?

Finally, calculate the F-Statistic or F-obtained?
Answer #1

#

ANOVA summary table

source Sum of Square df Mean Square F
Between groups 1,059 18 58.83 2.654
Within groups 3,702 167 22.168
Total 4,761 185

#Mean Squared Between =Between Sum of Square/df=1059/18=58.83

#Mean Squared Within=Within Sum of Square/df=37029/167=22.168

#F-Statistics=Mean Squared Between/Mean Squared Within

=58.33/22.168

=2.654